OpenStudy (anikate):

How do I factorize:x^2-5x+3

OpenStudy (lgbasallote):

does the question specifically asks you to factorize?

OpenStudy (anikate):

I need to find x intercepts so I need to know how to factorize it.

OpenStudy (lgbasallote):

since the last term is +3 the only possible factors are: (x+1)(x+3)<--produces 4x as middle term AND (x-1)(x-3)<--produces -4x as middle term therefore..it's not factorable...sadly..you have to use quadratic formula..

OpenStudy (anikate):

is that: \[(-b \pm \sqrt{b ^{2}-4ac})\div 2a\]

OpenStudy (lgbasallote):

yup that's the one!

OpenStudy (anikate):

there are 2 x intercepts in the answer

OpenStudy (lgbasallote):

since it's a quadratic equation..you expect that ;)

OpenStudy (anikate):

can you plz show me?

OpenStudy (anikate):

I can wait for a lon g time dont wry I'll be here

OpenStudy (anonymous):

the quadratic equation it gives u 2 x intercept depending on parabola or hyperbola curve

OpenStudy (anikate):

can u plz show how?

OpenStudy (lgbasallote):

your equation is x^2 - 5x + 3 a = 1 b = -5 c = 3 \[x = \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\] \[x = \frac{-(-5) \pm \sqrt{(-5)^2 - 4(1)(3)}}{2(1)}\] now do you still need help?

OpenStudy (anonymous):

that was easy

OpenStudy (anikate):

yes how did I get the x intercepts yet?

OpenStudy (lgbasallote):

the x intercept would be the value of that x

OpenStudy (lgbasallote):

you know how to solve that x right there right?

OpenStudy (anikate):

yes but I dont see how u got thst yet

OpenStudy (anikate):

then how would I get 2 x intercepts

OpenStudy (lgbasallote):

hmm i see your point let us solve it then... \[x = \frac{5 \pm \sqrt{25 - 12}}{2}\] \[x = \frac{5 \pm \sqrt{13}}{2}\] now do you see that \(\pm\) sign?

OpenStudy (anikate):

yup

OpenStudy (lgbasallote):

\(\pm\) means plus AND minus therefore \[x = \frac{5\pm \sqrt{13}}{2}\] this means \[x = \frac{5 + \sqrt{13}}{2}\] AND \[x = \frac{5 - \sqrt{13}}{2}\] so there are your TWO values